A 4-cylinder four stroke petrol engine develops 18.5 kW at 1500 r.p.m. The mean effective pressure is 5.5 bar. Calculate the stroke (in mm) the engine cylinder, if the length of stroke is 1.5 times the bore. [Write the answer with two decimal places but do not write the unit]


1
Expert's answer
2021-11-30T01:50:02-0500


i.P "=\frac{P_{i}×L×A×N×n}{60×z}"


Where i.P "=" Engine indicated power

"P_i=" Indicated mean effective pressure in pascals

"L=" Stroke length

"A=" Cylinder Area

"n=" Number of Cylinders

"z=2(" for 4 stroke engines")"

"60\to" Conversion of engine speed round per minute to round per second.


"\therefore18.5×1000=\frac{5.5×10^{5}(LA)×1500×4}{60×2}"


"LA=6.727\u00d710^{-4}m^3"


Let D "=" Diameter


"L=1.25D" and "A=\frac{\pi}{4}D^2"


"LA=(1.25D)(\frac{\pi}{4}D^2)=6.727×10^{-4}"


"D^3=6.852\u00d710^{-4}"

"D=8.816\u00d710^{-2}"


"L=1.5\u00d78.816\u00d710^{-2}"

"=0.1322m"

"=132.20"